# DAVOS: An Autonomous Vehicle Operating System in the Vehicle Computing Era Technical Report: CAR-TR-2025-009

### v1

@@ I Introduction
This separation leads to autonomy workloads and data-driven services coexisting through loosely coupled interfaces rather than sharing a common system substrate. As a result, resource management, data movement, and scheduling decisions are made independently across stacks, limiting system-level efficiency, extensibility, and coordination. These limitations motivate the need for an autonomous vehicle operating system designed explicitly for the vehicle computing era. Such a system should support predictable and safe execution for real-time driving workloads while also enabling extensible and secure data-driven services, bringing communication, computation, storage, and scheduling together within a unified system architecture.
@@ II Motivation
Figure 1: DAVOS architecture and vehicle deployment. In the figure, ADAS stands fors Advanced Driver Assistance Systems, AD denotes Autonomous Driving, RTOS stands for Real-Time Operating System, and ECUs represent Electronic Control Units. (a) DAVOS internal architecture, illustrating the data flow from hardware to applications and the separation between real-time ADAS/AD workloads and data-centric vehicle services. (b) Deployment of DAVOS on the vehicle central computer, interfacing with zonal gateways to support both driving and data services.
@@ II-B Limitations of Existing Vehicle Operating Systems
Current vehicle software stacks are built on a diverse set of operating systems and runtimes, each optimized for a specific domain, resulting in fragmented system support for autonomous and data-driven vehicle workloads. Infotainment and application-level services are commonly deployed on platforms such as Ford SYNC [19] and Android Automotive OS [26], which emphasize user interaction, application ecosystems, and connectivity. While these systems provide rich application support, they are not designed to offer predictable execution or timing awareness required by perception and control pipelines. As a result, they are typically isolated from driving critical workloads.
@@ V-A Problem Statement
At the same time, contemporary vehicle architectures divide computation across high-performance central compute units and smaller real-time domain controllers [6]. This produces two execution paths: a hard real-time bypass responsible for collision-avoidance safety [61], and a soft real-time main autonomy stack responsible for high-level decision-making. The main stack therefore does not need to meet deadlines with strict determinism, but it must continuously maintain sufficient quality in perception, prediction, and planning outputs to ensure stable and safe driving behavior [59].
@@ VI-B Architecture Overview
For each detected object, CRI computes three independent risk components based entirely on the CRI paper’s formulation: (1) Orientation risk, derived from the relative orientation angle between the ego and the object following Montewka et al.’s angular collision model. (2) Directional velocity risk, encapsulating longitudinal and lateral Time-to-Collision (TTC) using an approach-check mechanism: objects whose relative velocity indicates receding movement contribute zero risk, while approaching objects produce exponentially decaying TTC-based risk factors. (3) Speed-limit risk, modeled using a logistic function incorporating ego speed, road speed limits, and lane count following Kutela et al.
@@ VII-B Architecture Overview
INSIGHT™: Intelligent System for Incident Gathering, Handling, and Tracing (Tape). INSIGHT functions as the crash-aware forensic extension of AVS, responsible for safeguarding mission-critical evidence during and after abnormal driving events. Built upon the AVS ingestion services, it continuously monitors vehicle dynamics and sensor feedback to detect potential crash events in real time. Upon a confirmed incident, INSIGHT locks a pre- and post-crash time window from the SSD buffer and performs a write-once archival into the tape subsystem, ensuring tamper-proof preservation of all related data. Beyond storage, INSIGHT supports synchronized multi-sensor alignment and scene reconstruction, enabling precise post-event analysis for diagnostics, liability investigation, and safety model improvement.
@@ VIII-B Architecture Overview
Hardware VPIs. The Hardware VPI abstracts the physical components of the vehicle, providing a unified interface for sensors, actuators, controllers, communication devices, energy systems, and infotainment modules. It hides hardware differences across manufacturers and enables direct access to sensor configuration, actuator control, ECU management, and power operations. Sub-components include Sensor VPI for enumeration, calibration, and configuration; Actuator VPI for motion control; ECU VPI for subsystem coordination; Communication VPI for cellular, Wi-Fi, and V2X connectivity; Energy VPI for battery charging and power output; and Infotainment VPI for display and audio configuration. Together they enable standardized, hardware-agnostic access to the vehicle’s embedded devices.
@@ IX-B Threat Model
    Compromise user privacy by accessing sensitive personal data generated or stored within the vehicle
@@ IX-C Architecture Overview
Figure 7: PaCC system architecture. Green blocks represent core PaCC functions, red blocks represent algorithms run on AVS hardware, and blue blocks represent processing happens on PaCC Hardware.
@@
On the application side, every untrusted third party application must first be approved by the user and scanned by the system security services. Once approved, the system provisions a dedicated confidential execution container for the application. All computation of the application must take place inside this trusted execution environment, where memory, execution state, and internal storage remain isolated from the host.
@@
Through this design, PaCC ensures that even if the host system is compromised, all data stored in AVS remains encrypted, and all data processed by each application remains isolated and privacy filtered. PaCC protects the privacy of vehicle users while allowing approved third party services to access only the minimum necessary information. This architecture provides strong security at the storage level, execution level, and application level.

### v2

@@ I Introduction
This separation leads to autonomous workloads and data-driven services coexisting through loosely coupled interfaces rather than sharing a common system substrate. As a result, resource management, data movement, and scheduling decisions are made independently across stacks, limiting system-level efficiency, extensibility, and coordination. These limitations motivate the need for an autonomous vehicle operating system designed explicitly for the vehicle computing era. Such a system should support predictable and safe execution for real-time driving workloads while also enabling extensible and secure data-driven services, bringing communication, computation, storage, and scheduling together within a unified system architecture.
@@ II Motivation
Figure 1: DAVOS architecture and vehicle deployment. In the figure, ADAS stands for Advanced Driver Assistance Systems, AD denotes Autonomous Driving, RTOS stands for Real-Time Operating System, and ECUs represent Electronic Control Units. (a) DAVOS internal architecture, illustrating the data flow from hardware to applications and the separation between real-time ADAS/AD workloads and data-centric vehicle services. (b) Deployment of DAVOS on the vehicle central computer, interfacing with zonal gateways to support both driving and data services.
@@ II-B Limitations of Existing Vehicle Operating Systems
Current vehicle software stacks are built on a diverse set of operating systems and runtimes, each optimized for a specific domain, resulting in fragmented system support for autonomous and data-driven vehicle workloads. Infotainment and application-level services are commonly deployed on platforms such as Ford SYNC [19] and Android Automotive OS [26], which emphasize user interaction, application ecosystems, and connectivity. While these systems provide rich application support, they are not designed to offer the predictable execution or timing awareness required by perception and control pipelines. As a result, they are typically isolated from driving critical workloads.
@@ V-A Problem Statement
At the same time, contemporary vehicle architectures divide computation across high-performance central compute units and smaller real-time domain controllers [6]. This produces two execution paths: a hard real-time bypass responsible for collision-avoidance safety [61], and a soft real-time main autonomy stack responsible for high-level decision-making. The main stack, therefore, does not need to meet deadlines with strict determinism, but it must continuously maintain sufficient quality in perception, prediction, and planning outputs to ensure stable and safe driving behavior [59].
@@ VI-B Architecture Overview
For each detected object, CRI computes three independent risk components based entirely on the CRI paper’s formulation: (1) Orientation risk, derived from the relative orientation angle between the ego and the object following Montewka et al.’s angular collision model. (2) Directional velocity risk, encapsulating longitudinal and lateral Time-to-Collision (TTC) using an approach-check mechanism: objects whose relative velocity indicates receding movement contribute zero risk, while approaching objects produce decaying TTC-based risk factors exponentially. (3) Speed-limit risk, modeled using a logistic function incorporating ego speed, road speed limits, and lane count following Kutela et al.
@@ VII-B Architecture Overview
INSIGHT™: Intelligent System for Incident Gathering, Handling, and Tracing. INSIGHT functions as the crash-aware forensic extension of AVS, responsible for safeguarding mission-critical evidence during and after abnormal driving events. Built upon the AVS ingestion services, it continuously monitors vehicle dynamics and sensor feedback to detect potential crash events in real time. Upon a confirmed incident, INSIGHT locks a pre- and post-crash time window from the SSD buffer and performs a write-once archival into the subsystem, ensuring tamper-proof preservation of all related data. Beyond storage, INSIGHT supports synchronized multi-sensor alignment and scene reconstruction, enabling precise post-event analysis for diagnostics, liability investigation, and safety model improvement.
@@ VIII-B Architecture Overview
Hardware VPIs. The Hardware VPI abstracts the physical components of the vehicle, providing a unified interface for sensors, actuators, controllers, communication devices, energy systems, and infotainment modules. It hides hardware differences across manufacturers and enables direct access to sensor configuration, actuator control, ECU management, and power operations. Sub-components include Sensor VPI for enumeration, calibration, and configuration; Actuator VPI for motion control; ECU VPI for subsystem coordination; Communication VPI for cellular, Wi-Fi, and V2X connectivity; Energy VPI for battery charging and power output; and Infotainment VPI for display and audio configuration. Together, they enable standardized, hardware-agnostic access to the vehicle’s embedded devices.
@@ IX-B Threat Model
    Compromises user privacy by accessing sensitive personal data generated or stored within the vehicle
@@ IX-C Architecture Overview
Figure 7: PaCC system architecture. Green blocks represent core PaCC functions, red blocks represent algorithms run on AVS hardware, and blue blocks represent processing that happens on PaCC Hardware.
@@
On the application side, every untrusted third-party application must first be approved by the user and scanned by the system security services. Once approved, the system provisions a dedicated confidential execution container for the application. All computation of the application must take place inside this trusted execution environment, where memory, execution state, and internal storage remain isolated from the host.
@@
Through this design, PaCC ensures that even if the host system is compromised, all data stored in AVS remains encrypted, and all data processed by each application remains isolated and privacy filtered. PaCC protects the privacy of vehicle users while allowing approved third-party services to access only the minimum necessary information. This architecture provides strong security at the storage level, execution level, and application level.